# Feature Engineering is Not Dead: Reviving Classical Machine Learning with Entropy, HOG, and LBP Feature Fusion for Image Classification

### v1

@@ Abstract
Index Terms:
Permutation Entropy, Image Classification, Feature Fusion, Classical Machine Learning, Interpretable Models, Histogram of Oriented Gradients (HOG), Local Binary Patterns (LBP), Support Vector Machines (SVM), Benchmark Datasets (e.g., Fashion-MNIST, CIFAR-10).
@@ I Introduction
Permutation entropy (PE), introduced by Bandt and Pompe [1], is a symbolic time-series complexity measure that captures the diversity of ordinal patterns in a signal. Unlike conventional entropy metrics such as Shannon or Rényi entropy [2, 3], which are derived from amplitude-based probability distributions, PE encodes local temporal dynamics by analyzing the relative ordering of values within sliding windows. Due to its robustness to noise, computational simplicity and model-free formulation, PE has found widespread application in physiological signal analysis, including EEG-based seizure detection [4], heart rate variability [5], and sleep stage classification [6].
Despite its success in 1D temporal domains, PE has rarely been explored in image classification. A few works have used PE in medical imaging scenarios—such as patch-wise entropy for CT [7] and fMRI [8] analysis—but these have been limited in scope, focusing on narrow domains or local feature statistics. To the best of our knowledge, PE has never been employed as a primary, global feature descriptor for natural image classification.
@@
In this work, we present a novel, interpretable, and computationally lightweight approach that uses PE for full-image classification. The PE features are combined with two complementary image descriptors, the Histogram of Oriented Gradients (HOG) [14] and the Local Binary Patterns (LBP) [15]. HOG captures shape and edge structure of an image, useful for detecting object outlines and silhouettes, whereas LBP captures micro-texture such as spots, flat areas, edges and corners, and good at capturing fine repetitive patterns. We combine PE, HOG and LBP features to form a compact, handcrafted feature vector of n-dimensions. We use these handcrafted feature vectors and employ Support Vector Machines (SVM) [16], a classical machine learning method, for image classification work across multiple benchmark datasets such as Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10. In particular, we achieve up to 91.15% test accuracy on Fashion-MNIST benchmark dataset. All previous test accuracy results [17] in the Fashion-MNIST benchmark dataset, that used scikit-learn library and evaluated 129 classifiers with various parameter configurations (excluding deep learning methods), reported a maximum accuracy of 89.7% using SVM with parameters C=10, kernel=poly [17]. In contrast, our method achieves an accuracy of 91.15% for the first time using purely classical machine learning techniques applied on hand-crafted features extracted from the Fashion-MNIST dataset. We also apply this method on other benchmark datasets (KMNIST, EMNIST, and CIFAR-10) and show competitive performance.
The rest of the paper is organized as follows: In Section II, we present a detailed review of permutation entropy; Section III discusses the classical image descriptors HOG and LBP employed in this work; Section IV describes our feature-extraction pipeline in detail; Section V briefly introduces classical SVM-based machine learning; Section VI explains our rationale for selecting the benchmark datasets and provides their descriptions; Section VII reports the results obtained using our novel feature-extraction pipeline and compares them with previous studies; finally, Section VIII concludes the paper and outlines future implications.
@@ III-A Histogram of Oriented Gradients (HOG)
Figure 2: Visualization of HOG from an image: (a) reference image (source: scikit-image); (b) corresponding HOG; (c) Schematic visualization of 9 HOG Orientation Bins (0°–180°, unsigned; arrows = gradient directions). Note that a gradient at 90° (arrow pointing straight up) comes from a horizontal edge, while a gradient at 0°/180° (arrow along the x-axis) comes from a vertical edge; (d) plot of global HOG orientation histogram (9 orientation bins, 0°–180°) in which each bar aggregates the total gradient magnitude for orientations falling into its 20° bin.
@@
HOG features are translation-invariant at the cell level and capture edge and contour information that is highly discriminative for structured objects.
@@ III-B Local Binary Pattern (LBP)
Figure 3: Visualization of LBP from an image: (a) reference image (source: scikit-image); (b) corresponding LBP.
Combined Use in Feature Fusion: When used together, HOG and LBP provide complementary information: HOG captures global structure and shape, while LBP encodes fine-scale local texture. Note that although HOG is more of an edge/shape descriptor, but its dense‐grid histogram approach can incidentally pick up some texture cues (Fig. 2). In contrasts, LBP is explicitly a texture descriptor (Fig. 3). Feature fusion of HOG and LBP has been shown to improve classification performance in various tasks, including face recognition, scene classification, and biomedical imaging [19].
@@ IV-A Feature Extraction Process and Dimensions
In this work, each image from standard benchmark datasets (such as Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10) is transformed into a handcrafted feature vector composed of entropy-based, correlation-based, shape and texture-based image descriptors. Although dimensionality may vary slightly between datasets due to image size or content, the core extraction mechanisms remain consistent (Fig. 4). In the following, we provide a general overview of the feature types and the methods used for their extraction.
@@
Diagonal PE: We simplify the diagonal feature extraction as follows. Starting with the main diagonal (top-left to bottom-right), we also take the $K$ diagonals above it and the $K$ diagonals below it—so in total $12K+$ 1 diagonals. For each diagonal line, we read its pixel values in order to form a one-dimensional sequence, then compute permutation entropy on that sequence and again on its reversed sequence. By taking the geometric mean of these two entropy values, we collapse each diagonal into a single feature . This reflects how complex the pixel arrangement is along that slanted direction.
@@
$PE_{\text{diag}}(k)=\frac{PE(d_{k})+PE(d_{k}^{\text{rev}})}{2}$
Anti-Diagonal PE: We handle anti-diagonals in exactly the same way. An anti-diagonal is simply a line of pixels running from the top-right corner down to the bottom-left corner (and any lines parallel to it). Concretely, we take the main anti-diagonal (top-right to bottom-left) and its $K$ offsets on either side, thus again yielding $12K+$ 1 lines—read each line’s pixel values into a one-dimensional sequence, compute permutation entropy on that sequence and on its reverse, then combine them via the geometric mean. This produces one scalar feature per anti-diagonal that, like the diagonal features, captures the complexity of pixel arrangements along that slanted orientation.
@@
$PE_{\text{anti-diag}}(k)=\frac{PE(ad_{k})+PE(ad_{k}^{\text{rev}})}{2}$
Patch-wise PE: The image is divided into overlapping patches (e.g., $4\times 4$ with stride 2), and the PE is calculated for each patch:
@@
Figure 4: (a) Schematic overview of the proposed workflow pipeline. For each input image we extract permutation-entropy (PE) based features—computed along rows, columns, the diagonal and anti-diagonals, and on overlapping local patches. We combine them with row-/column-correlation based features, histogram of oriented gradients (HOG) and local binary pattern (LBP) image descriptors. The four feature sets are concatenated into a single feature vector and classified with a support vector machine (SVM). (b) Detailed illustration of the PE-based feature-extraction steps on a (6 × 6) toy image. PE is first evaluated for every row and column, followed by the diagonal (top-left → bottom-right) and anti-diagonal (top-right → bottom-left) directions. Note that some off-centre diagonals and anti-diagonals (towards corner) PEs are not computed because their reduced length may yield unreliable PE features or statistics. Finally, we compute patch-wise PE with (4 × 4) patches and with a stride length of 2 pixels. See Section IV for a complete description of the feature-extraction pipeline.
@@ 6 $a_{61}$ $a_{62}$ $a_{63}$ $a_{64}$ $a_{65}$ $a_{66}$
[…]
Let us now move to Diagonal-wise PE with $K=2$. For each offset $k=-2,-1,0,1,2$ define the diagonal
[…]
@@
Similarly, for each $k=-2,-1,0,1,2$ define the anti‐diagonal
[…]
@@
$PE_{\mathrm{anti}}^{GM}(k)=\sqrt{\,PE(ad_{k})\,\times\,PE(ad_{k}^{\mathrm{rev}})\,}\,,\quad k=-2,\ldots,2,\\ \$
@@ V Support Vector Machine (SVM) for image classification
[…]
@@ VII-A Results on Fashion-MNIST Dataset
Row-wise PE: 28 rows → 28 features; Column-wise PE: 28 columns → 28 features; Row Correlation: 27 adjacent row pairs → 27 features; Column Correlation: 27 adjacent column pairs → 27 features; Diagonal PE: 21 diagonals ($k=-10$ to $10$) → 21 features; Anti-diagonal PE: 21 diagonals on horizontally flipped image → 21 features; Patch-wise PE: 169 overlapping patches of size $4\times 4$ with stride 2 → 169 features; HOG: 441 features using $4\times 4$ pixel cells and 9 orientation bins; LBP: 18-bin histogram with $P=16$, $R=2$ (uniform pattern) → 18 features.
TABLE II: Breakdown of Handcrafted Feature Types Used in Fashion-MNIST Dataset
Feature Type Number of Features
Row-wise PE 28
Column-wise PE 28
Diagonal PE 21
Anti-diagonal PE 21
Patch-wise PE 169
Row Correlation 27
Column Correlation 27
HOG 441
LBP 18
@@
This 780-dimensional feature vector (Table II) is standardized and used as input to the classification models. Similar extraction procedures are applied to other datasets (KMNIST, EMNIST, CIFAR-10), with minor adjustments to patch count or image resolution as necessary. The resulting feature matrices had shapes $(60000,780)$ for training and $(10000,780)$ for testing. An SVM classifier with RBF kernel was tuned using 3-fold cross-validation over a grid of 12 $(C,\gamma)$ combinations, totaling 36 fits. The best hyperparameters obtained were $C=10$ and $\gamma=0.001$.The best cross-validation accuracy achieved during tuning was 90.44%, and the final test accuracy using the best SVM model was 91.15%. Table III presents a comparison of our results with previously reported SVM-based approaches on the Fashion-MNIST dataset that did not employ deep learning or handcrafted descriptors of comparable complexity.
@@
Ours (PE + Corr + HOG + LBP) 780 handcrafted features 91.15%

### v2

@@ I Introduction
Permutation entropy (PE), introduced by Bandt and Pompe [1], is a symbolic time-series complexity measure that captures the diversity of ordinal patterns in a signal. Unlike conventional entropy metrics such as Shannon or Rényi entropy [2, 3], which are derived from amplitude-based probability distributions, PE encodes local temporal dynamics by analyzing the relative ordering of values within sliding windows. Due to its robustness to noise, computational simplicity and model-free formulation, PE has found widespread application in physiological signal analysis, including EEG-based seizure detection [4], heart rate variability [5], and sleep stage classification [6], , and fMRI  [7] analysis.
Despite its success in 1D temporal domains, permutation entropy (PE) remains largely unexplored in natural image classification. A few previous works have used entropy-related ideas in medical imaging—for example, entropy-guided patch sampling within CT liver segmentation pipelines [8]—but these approaches are task- and modality-specific and do not treat PE as a general-purpose image descriptor. Beyond medical image analysis, entropy has a long-standing role in classical image processing. Early work used entropy principles for fundamental operations such as gray-level thresholding and segmentation [9]. More recently, several two-dimensional ordinal/entropy variants have been proposed specifically for texture characterization and discrimination  [10, 11, 12]. However, to the best of our knowledge, PE has not been employed as a primary, global handcrafted feature for natural image classification within a unified multi-orientation, multi-scale design.
@@
In this work, we present a novel, interpretable, and computationally lightweight approach that uses PE for full-image classification. The PE features are combined with two complementary image descriptors, the Histogram of Oriented Gradients (HOG) [18] and the Local Binary Patterns (LBP) [19]. HOG captures shape and edge structure of an image, useful for detecting object outlines and silhouettes, whereas LBP captures micro-texture such as spots, flat areas, edges and corners, and good at capturing fine repetitive patterns. We combine PE, HOG and LBP features to form a compact, handcrafted feature vector of n-dimensions. Let us clarify in what sense the handcrafted features are interpretable. By interpretable, we mean that each feature group has a clear, known semantic meaning. HOG captures oriented edges and shape, LBP encodes local texture patterns, and PE measures local complexity and irregularity. Unlike CNNs, where features are learned automatically and often require post-hoc tools (e.g., Grad-CAM, SHAP) to explain decisions, our descriptors are predefined and semantically grounded, allowing us to directly identify what image properties the classifier relies upon.
We use these handcrafted feature vectors and employ Support Vector Machines (SVM) [20], a classical machine learning method, for image classification work across multiple benchmark datasets such as Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10. In particular, we achieve up to 91.23% test accuracy on Fashion-MNIST benchmark dataset. All previous test accuracy results [21] in the Fashion-MNIST benchmark dataset, that used scikit-learn library and evaluated 129 classifiers with various parameter configurations (excluding deep learning methods), reported a maximum accuracy of 89.7% using SVM with parameters C=10, kernel=poly [21]. In contrast, our method achieves an accuracy of 91.23% for the first time using purely classical machine learning techniques applied on hand-crafted features extracted from the Fashion-MNIST dataset. We also apply this method on other benchmark datasets (KMNIST, EMNIST, and CIFAR-10) and show competitive performance.
The rest of the paper is organized as follows: In Section II, we present a detailed review of permutation entropy; Section III discusses the classical image descriptors HOG and LBP employed in this work; Section IV describes our feature-extraction pipeline in detail; Section V briefly introduces classical SVM-based machine learning; Section VI explains our rationale for selecting the benchmark datasets and provides their descriptions; Section VII reports the results obtained using our novel feature-extraction pipeline and compares them with previous studies; finally, Section IX concludes the paper and outlines future implications.
@@ III-A Histogram of Oriented Gradients (HOG)
Figure 2: Visualization of HOG from an image: (a) reference image (source: scikit-image); (b) corresponding HOG that nicely captures shape, outline and edges of the image; (c) Schematic visualization of 9 HOG Orientation Bins (0°–180°, unsigned; arrows = gradient directions). Note that a gradient at 90° (arrow pointing straight up) comes from a horizontal edge, while a gradient at 0°/180° (arrow along the x-axis) comes from a vertical edge; (d) plot of global HOG orientation histogram (9 orientation bins, 0°–180°) in which each bar aggregates the total gradient magnitude for orientations falling into its 20° bin.
@@
HOG features are translation-invariant at the cell level and capture edge and contour information that is highly discriminative for structured objects. In addition, HOG remains a widely used handcrafted descriptor in applied vision pipelines due to its robustness in capturing gradient/edge structure. For example, Sun et al. used a vision plus SVM pipeline for train classification by converting weigh-in-motion (WIM) time series into images and extracting visual features for SVM-based categorization  [35]. In agricultural machine vision, Xu et al. proposed a two-stage strawberry detection method that combines HSV-based candidate localization with an HOG+SVM classifier to handle slightly overlapping fruit  [36]. The studies as above motivate HOG as a strong, task-relevant shape/edge baseline, which we include as one component in our hybrid feature stack alongside permutation-entropy complexity cues and LBP texture descriptors.
@@ III-B Local Binary Pattern (LBP)
Figure 3: Visualization of LBP from an image: (a) reference image (source: scikit-image); (b) corresponding LBP which is good at encoding fine-scaled local texture.
Combined Use in Feature Fusion: When used together, HOG and LBP provide complementary information: HOG captures global structure and shape, while LBP encodes fine-scaled local texture. Note that although HOG is more of an edge/shape descriptor, but its dense‐grid histogram approach can incidentally pick up some texture cues (Fig. 2). In contrasts, LBP is explicitly a texture descriptor (Fig. 3). Feature fusion of HOG and LBP has been shown to improve classification performance in various tasks, including face recognition, scene classification, and biomedical imaging [25].
@@ IV-A Feature Extraction Process and Dimensions
In this work, each image from standard benchmark datasets (such as Fashion-MNIST, KMNIST, EMNIST, and CIFAR-10) is transformed into a handcrafted feature vector composed of entropy-based, correlation-based, shape and texture-based image descriptors. Although dimensionality may vary slightly between datasets due to image size or content, the core extraction mechanisms remain the same (Fig. 4). In the following, we provide a general overview of the feature types and the methods used for their extraction.
@@
Diagonal PE: We simplify the diagonal feature extraction as follows. Starting with the main diagonal (top-left to bottom-right), we also take the $K$ diagonals above it and the $K$ diagonals below it—so in total $2K+1$ diagonals (Fig. 4b). For each diagonal line, we read its pixel values in order to form a one-dimensional sequence, then compute permutation entropy on that sequence and again on its reversed sequence. By taking the geometric mean of these two entropy values, we collapse each diagonal into a single feature . This reflects how complex the pixel arrangement is along that slanted direction.
@@
$PE_{\text{diag}}(k)=\sqrt{PE(d_{k})\cdot PE(d_{k}^{\text{rev}})}$
Anti-Diagonal PE: We handle anti-diagonals in exactly the same way. An anti-diagonal is simply a line of pixels running from the top-right corner down to the bottom-left corner (and any lines parallel to it; Fig. 4b). Concretely, we take the main anti-diagonal (top-right to bottom-left) and its $K$ offsets on either side, thus again yielding $2K+1$ lines—read each line’s pixel values into a one-dimensional sequence, compute permutation entropy on that sequence and on its reverse, then combine them via the geometric mean. This produces one scalar feature per anti-diagonal that, like the diagonal features, captures the complexity of pixel arrangements along that slanted orientation.
@@
$PE_{\text{anti-diag}}(k)=\sqrt{PE(ad_{k})\cdot PE(ad_{k}^{\text{rev}})}$
Patch-wise PE: The image is divided into overlapping patches (e.g., $4\times 4$ with stride 2; Fig. 4b), and the PE is calculated for each patch:
@@
Figure 4: (a) Schematic overview of the proposed workflow pipeline. For each input image we extract permutation-entropy (PE) based features—computed along rows, columns, the diagonal and anti-diagonals, and on overlapping local patches. We combine them with row-/column-correlation based features, histogram of oriented gradients (HOG) and local binary pattern (LBP) image descriptors. The four feature sets are concatenated into a single feature vector and classified with a support vector machine (SVM). (b) Detailed illustration of the PE-based feature-extraction steps on a (6 × 6) toy image. PE is first evaluated for every row and column, followed by the diagonal (top-left → bottom-right) and anti-diagonal (top-right → bottom-left) directions. Note that some off-centre diagonals and anti-diagonals (towards corner) PEs are not computed because their reduced length may yield unreliable PE features or statistics. Finally, we compute patch-wise PE with (2 × 2) patches and with a stride length of 2 pixels. See Section IV for a complete description of the feature-extraction pipeline.
@@ 6 $a_{61}$ $a_{62}$ $a_{63}$ $a_{64}$ $a_{65}$ $a_{66}$
yielding six column‐wise features $\{PE_{\mathrm{col},\,1}^{GM},\,PE_{\mathrm{col},\,2}^{GM},\,\dots,\,PE_{\mathrm{col},\,6}^{GM}\}$.
Let us now move to Diagonal-wise PE with $K=2$. For each offset k=2,1,0,1,2k=-2,-1,0,1,2 define the diagonal
$d_{k}$ $=\bigl(a_{p,\;p+k}\;\bigm|\;1\leq p\leq 6,\;1\leq p+k\leq 6\bigr),$
@@
Similarly, for each k=2,1,0,1,2k=-2,-1,0,1,2 define the anti‐diagonal
$ad_{k}=\bigl(a_{p,\;7-p+k}\;\bigm|\;1\leq p\leq 6,\;1\leq 7-p+k\leq 6\bigr),$
@@
$PE_{\mathrm{anti}}^{GM}(k)=\sqrt{\,PE(ad_{k})\,\times\,PE(ad_{k}^{\mathrm{rev}})\,}\,,\quad k=-2,\ldots,2,\\$ \\
@@ V Support Vector Machine (SVM) for image classification
$\max\bigl(0,\,1-y_{i}(\mathbf{w}^{T}\phi(\mathbf{x}_{i})+b)\bigr)$ — The hinge loss for example $i$: zero if correctly classified with margin at least 1, and growing linearly when the functional margin $y_{i}(\mathbf{w}^{T}\phi(\mathbf{x}_{i})+b)$ falls below 1.
@@ VII-A Results on Fashion-MNIST Dataset
Row-wise PE: 28 rows → 28 features; Column-wise PE: 28 columns → 28 features; Row Correlation: 27 adjacent row pairs → 27 features; Column Correlation: 27 adjacent column pairs → 27 features; Diagonal PE: 21 diagonals (k=10k=-10 to $10$) → 21 features; Anti-diagonal PE: 21 diagonals on horizontally flipped image → 21 features; Patch-wise PE: 169 overlapping patches of size $4\times 4$ with stride 2 → 169 features; HOG: 441 features using $4\times 4$ pixel cells and 9 orientation bins; LBP: 18-bin histogram with $P=16$, $R=2$ (uniform pattern) → 18 features.
TABLE II: Breakdown of handcrafted feature types used in Fashion-MNIST.
Feature Type (Count rationale) Features
Row-wise PE (28 rows) 28
Column-wise PE (28 columns) 28
Row Correlation (27 adjacent row pairs) 27
Column Correlation (27 adjacent column pairs) 27
Diagonal PE ($k=-10,\ldots,10$) 21
Anti-diagonal PE (flip + $k=-10,\ldots,10$) 21
Patch-wise PE ($4{\times}4$, stride 2 $\Rightarrow 13{\times}13$) 169
HOG ($4{\times}4$ cells, 9 bins $\Rightarrow 7{\times}7{\times}9$) 441
LBP ($P=16$, $R=2$ $\Rightarrow P{+}2$ bins) 18
@@
This 780-dimensional feature vector (Table II) is standardized and used as input to the classification models. Similar extraction procedures are applied to other datasets (KMNIST, EMNIST, CIFAR-10), with minor adjustments to patch count or image resolution as necessary. The resulting feature matrices had shapes $(60000,780)$ for training and $(10000,780)$ for testing. An SVM classifier with RBF kernel was tuned using 3-fold cross-validation over a grid of 12 $(C,\gamma)$ combinations, totaling 36 fits. The best hyperparameters obtained were $C=10$ and $\gamma=0.001$.The best cross-validation accuracy achieved during tuning was 90.44%, and the final test accuracy using the best SVM model was 91.23%. Table III presents a comparison of our results with previously reported SVM-based approaches on the Fashion-MNIST dataset that did not employ deep learning or handcrafted descriptors of comparable complexity.
@@
PE/Corr + HOG + LBP + SVM 780 handcrafted features 91.23%
@@
TABLE IV: Fashion-MNIST ablation (60k/10k split). Accuracy is reported using a fixed RBF-SVM configuration (C=10C{=}10, $\gamma{=}0.001$) for a controlled comparison across feature sets.
Feature set Dim. Acc. (%)
PE/Corr 321 87.31
HOG 441 89.28
LBP 18 55.34
HOG+LBP 459 89.90
PE/Corr+HOG 762 90.87
PE/Corr+LBP 339 87.74
Full (PE/Corr+HOG+LBP) 780 91.23
TABLE V: Sensitivity of the PE/Corr feature block (321-D) to the permutation entropy parameters: embedding dimension $m$ and delay $\tau$ on Fashion-MNIST (60k/10k split). Accuracy is reported using a fixed RBF-SVM configuration (C=10C{=}10, $\gamma{=}0.001$) for a controlled comparison. Note: $m$ is the PE embedding dimension (pattern length); the “2” indicates that the header cell spans two rows for formatting and is not $2{\times}m$.
$m$ $\tau$
1 2 3
3 86.88 87.29 87.06
4 87.31 86.91 86.92
5 86.96 86.69 86.80
Tables IV and V provide parameter sensitivity for the proposed design choices on Fashion-MNIST. The ablation in Table IV shows that the PE/Corr block alone is competitive (87.31%), while HOG contributes the strongest individual gain (89.28%); combining PE/Corr with HOG improves accuracy to 90.87%, and the full descriptor achieves the best performance (91.23%), indicating that ordinal-entropy/correlation statistics provide information complementary to gradient-based texture cues. To address sensitivity of permutation entropy to its core parameters, Table V varies the embedding dimension $m$ and delay $\tau$. Performance is stable across the tested range ($m\in{3,4,5}$ and $\tau\in{1,2,3}$), with a small spread of approximately 0.6 percentage points, suggesting that the PE/Corr representation is not overly sensitive to moderate changes in $(m,\tau)$ and supporting the robustness of the chosen default configuration.
For comparison of performance with previous tensor-based HOG studies  [22]  [23], we additionally evaluated HOG as a tensor of size $7\times 7\times 9$ ($_{\text{row}} ×\times blocks$$_{\text{col}}$ $\times$ orientations) within the same hybrid feature stack (Entropy + HOG + LBP). As shown in Table VI, naive tensor flattening yields essentially the same performance as standard vector-HOG (91.19% vs. 91.23%), consistent with both representations containing the same HOG coefficients. MPCA (ranks $(7,7,9)$) slightly reduces accuracy (90.57%), and a tensor-aware STM baseline (OVR, CP rank 8) achieves 72.94% accuracy; thus, these tensor-based baselines do not improve over the tuned RBF-SVM reference in our setting.
TABLE VI: Tensor-based HOG formulations under same hybrid feature stack on Fashion-MNIST.
Method HOG Representation / Tensor Method Accuracy
Baseline (Vector-HOG + SVM) Vector-HOG within 780 handcrafted features 91.23%
Tensor-HOG (flatten) + SVM Tensor-HOG ($7\times 7\times 9$) flattened (same coefficients) 91.09%
Tensor-HOG + MPCA + SVM MPCA ranks $(7,7,9)$ (core dim = 441) 90.57%
Tensor-aware classifier (STM-OVR) STM-OVR (CP rank = 8) on tensor-HOG; Entropy+LBP unchanged 72.94%